: Solve Ricatti’s DEq. dy/dx =-2-y-y2 , y1=2


1
Expert's answer
2021-05-07T09:46:08-0400

Given Differential equation: "\frac{dy}{dx}" = 2 - y - y2 , y(1) = 2


since it is a separable differential equation, therefore, separating it we get,

"\implies" "\frac{dy}{2 - y - y^2}" = "dx"

taking integral both side,

"\implies" "\int" "\frac{dy}{2 - y - y^2}" = "\int" "dx"


"\implies" "\int" "\frac{dy}{2 - y - y^2}" = "\int" "dx"


by using partial fraction decomposition we can write,


"\frac{dy}{2 - y - y^2}" = "dy" ( "\frac{-1}{3(y-1)}" + "\frac{1}{3(y+2)}" ) = "\frac{-1}{3(y-1)}dy" + "\frac{1}{3(y+2)}dy"


therefore,


"\implies" "\int"( "\frac{-1}{3(y-1)}dy" + "\frac{1}{3(y+2)}dy" ) = "\int" "dx"


"\implies" "\int" "\frac{-1}{3(y-1)}dy" + "\int" "\frac{1}{3(y+2)}dy" = "\int" "dx"


"\implies" "-\frac{1}{3}\ln \left(y-1\right)+\frac{1}{3}\ln \left(y+2\right)=t+C"


since given y(1) = 2


"-\frac{1}{3}\ln \left(2-1\right)+\frac{1}{3}\ln \left(2+2\right)=1+C"


"\implies" C = "\frac{2\ln \left(2\right)}{3}-1"

hence solution of given Ricatti’s DEq is given by,


"\implies" "-\frac{1}{3}\ln \left(y-1\right)+\frac{1}{3}\ln \left(y+2\right)=t+\frac{2\ln \left(2\right)}{3}-1"